OpenStudy (anonymous):

HELP PLEASE!!!! Find the area of a sector with an arc length of 60 in. and a radius of 15 in. A. 117.75 in.² B. 282.60 in.² C. 450 in.² D. 6,750 in.²

OpenStudy (jiteshmeghwal9):

Area of a sector=\(\Large{\frac{\theta}{360} \times \pi r^2}\)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

would the answer be A?

OpenStudy (anonymous):

because it says its wrong

OpenStudy (anonymous):

is 60 the theta

jimthompson5910 (jim_thompson5910):

60 is not the theta 60 is the distance around the curve. It's the length of the arc this shaded distance 

jimthompson5910 (jim_thompson5910):

if the radius is 15, then what is the total circumference?

OpenStudy (anonymous):

30?

jimthompson5910 (jim_thompson5910):

use the formula C = 2*pi*r

jimthompson5910 (jim_thompson5910):

you're close

OpenStudy (anonymous):

i got 94.24

OpenStudy (anonymous):

when i plugged it in

jimthompson5910 (jim_thompson5910):

yeah \[\Large 30\pi \approx 94.2477796\]

jimthompson5910 (jim_thompson5910):

I'm going to leave it as 30pi to keep things as exact as possible

OpenStudy (anonymous):

okay so what i do with that